I am having trouble solving the following problem. Please help.
seven thirds x = two thirds. It is a fraction equation.

Answer by wuwei96815(245)

If I understand the problem correctly it is:

7/3x = 2/3

First divide both sides by 7/3 to isolate x:
7/3x / 7/3 = 2/3 / 7/3
x = 2/3 / 7/3 = 2/3 times 3/7 = 2/7
x = 2/7

Substitute this answer into the original equation to see if it is correct.